anoumous1010:

To mix a punch, Lindsay begins with 3 liters of lemon-lime soda and adds 1,893 milliliters of cranberry juice. How many liters are in the punch bowl when the punch is completed

Elsa213:

Maybe 4.893 o:

anoumous1010:

okie '

Hero:

@kylieisweird101 can you please go back to your question?

Hero:

@anoumous1010 covert milliliters to liters. Remember milli = 1/1000

Hero:

\(1.893 \text{ milliliters} = 1.893\left(\dfrac{1}{1000}\right)\text{ liters}\)